# Energy-Efficient Split Learning for Fine-Tuning Large Language Models in Edge Networks

### v1

@@ I Introduction
Parameter efficient fine-tuning (PEFT) has been explored to reduce the number of trainable parameters for LLMs, which can enhance the performance of pre-trained LLMs for specific tasks in edge networks [3]. It only fine-tunes a small number of model parameters instead of all the model’s parameters while maintaining comparable performance with the full-model fine-tuning. In particular, as a popular PEFT method, low-rank adaptation (LoRA) keeps the pre-trained model weights fixed and introduces trainable low-rank matrices within each transformer layer. The LLM fine-tuning has demonstrated advanced performance, but many private institutions or companies have no sufficient privileges to access the database stored on edge users due to concerns about data privacy and security [4].
@@
The federated learning (FL) paradigm emerges as a promising solution to address these privacy concerns by enabling decentralized training of LLMs directly on multiple devices without sharing raw data, instead exchanging only model updates or gradients [5]. However, LoRA-based fine-tuning an LLM for a specific task requires enormous computational overhead and memory, hindering its execution on low-end edge devices [6]. For example, fine-tuning a T5-Large model with LoRA requires memory of 7.1 GB in a mini-batch training, while the dynamic random-access memory (DARM) of a Jetson Nano (a dedicated platform for edge AI) is only 4 GB [7].
@@
To address this issue, split learning (SL) has been introduced as an alternative approach that reduces the computational burden on edge devices while preserving data privacy [8]. In SL, the initial layers of a model are processed locally on the device, and only the intermediate activations (not the raw data) are sent to the server, which then completes the remaining layers of the model [9]. In [4], an SL-based fine-tuning approach is introduced, in which the edge devices fine-tune the initial layers of the LLM while edge servers execute the remaining layers. Similarly, a split LoRA-based fine-tuning framework in wireless networks is presented, where the embedding and task modules are executed on the devices, with the encoder kept on the server [10]. Most existing literature primarily focuses on static split strategies and does not consider the impact of server-side energy consumption.
@@
However, there are still a variety of challenges in how LLMs can be optimized in edge networks to achieve the best possible performance in both latency and energy efficiency. Firstly, it is suboptimal to use a static cut layer to split the model training process between mobile devices and the edge server, due to the device heterogeneity and channel dynamics in edge networks. Secondly, a static configuration for the server’s computational resource can lead to suboptimal energy usage, as its processing workload can vary significantly based on the cut layer. Therefore, an adaptive strategy that dynamically adjusts the cut layer and server’s computational resource based on real-time conditions, is necessary to optimize both latency and energy efficiency.
@@
In this letter, firstly, we propose an energy-efficient SL framework for fine-tuning LLMs, which aims to enhance the training and energy efficiency in edge networks. Secondly, we analyze the factors of the training delay and energy consumption in detail, and then formulate an optimization problem to minimize both the training delay and the server’s energy consumption. Thirdly, to solve this problem, we design an algorithm of Cut lAyer and computing Resource Decision, named CARD. It decomposes this problem into independent sub-problems, and then jointly determines the cut layer selection and server’s computation resource allocation. Finally, extensive simulations demonstrate the performance of the proposed approach through a physical platform with 5 NVIDIA Jetson edge devices. The main contributions of this work are summarized as follows:
@@ II-A LoRA-Based LLM Fine-Tuning
A typical LoRA-based LLM fine-tuning scenario is considered, in which multiple edge devices collaborate with an access point (AP) to fine-tune a shared pre-trained LLM. Each device trains only the initial layers of the LLM, reducing local computation and memory requirements. The AP with an edge server coordinates to complete the remaining training process with each device, as depicted in Fig. 1. This approach is ideal for resource-constrained environments, enabling collaborative fine-tuning while enhancing the training and energy efficiency.
Let $\textbf{W}\in\mathbb{R}^{P\times Q}$ denote a pre-trained weight matrix of the LLM, with its corresponding model update expressed as $\textbf{W}+\Delta\textbf{W}=\textbf{W}+\textbf{A}\textbf{B}$. Here, $\textbf{A}\in\mathbb{R}^{P\times Z}$ and $\textbf{B}\in\mathbb{R}^{Z\times Q}$ are a pair of decomposed low-rank matrices, and $Z\ll\min(P,Q)$, as shown in Fig. 2. In the LoRA-based fine-tuning process, the pair of decomposed low-rank matrices are used for adaptation, while the weight matrix W is frozen. The edge device is indexed by $m\in\mathcal{M}=\{1,2,...,M\}$, and its local dataset is denoted by $\mathcal{D}_{m}=\{\textbf{x}_{m,k},y_{m,k}\}_{k=1}^{D_{m}}$. Here, $\textbf{x}_{m,k}$ and $y_{m,k}$ are the $k$-th input data and its corresponding label, respectively. A device-side pre-trained model and its associated set of trainable LoRA adapters are denoted by $\textbf{W}_{m}^{D}$ and $\textbf{R}_{m}^{D}$, respectively. $\textbf{R}_{m}^{D}=\{\textbf{A}_{m}^{1},\textbf{B}_{m}^{1},\dots,\textbf{A}_{m}^{c_{m}},\textbf{B}_{m}^{c_{m}}\}$, where $\textbf{A}_{m}^{i}$ and $\textbf{B}_{m}^{i}$ are the pair of low-rank matrices of $i$-th LoRA adapter, and $c_{m}$ is the cut layer.
@@
For the edge server deployed in the AP, a server-side pre-trained model and its corresponding set of trainable LoRA adapters are denoted by $\textbf{W}_{m}^{S}$ and $\textbf{R}_{m}^{S}$, respectively. Similarly, $\textbf{R}_{m}^{S}=\{\textbf{A}_{m}^{c_{m}+1},\textbf{B}_{m}^{c_{m}+1},\dots,\textbf{A}_{m}^{I},\textbf{B}_{m}^{I}\}$, where $I$ is the total number of transformer layers in the LLM. The global pre-trained LLM is denoted by $\textbf{W}=\{\textbf{W}_{m}^{D};\textbf{W}_{m}^{S}\},~{}\forall m\in\mathcal{M}$. The objective of split LoRA-based fine-tuning is to find the optimal LoRA adapter configuration that minimizes the global loss across all participating devices, i.e.,
@@ II-B Proposed SL Framework for fine-tuning LLMs
In this section, the proposed SL framework for fine-tuning LLMs is presented in detail. The basic idea is leveraging the SL method allowing devices to fine-tune the first layers of the LLM while the server handles the remaining layers. Before the initiation of fine-tuning, the parameters in LoRA adapters are initialized randomly. After that, the LLM is executed between devices and the server in consecutive training rounds until the optimal LoRA adapters are identified. The training round is indexed by $n\in\mathcal{N}=\{1,2,...,N\}$, and the detailed workflow of the proposed framework is as follows.
Stage 1 - LLM splitting. The edge server selects device $m$ to collectively execute LLM fine-tuning. Then, it splits the trainable LoRA adapters into device-side LoRA adapters $\textbf{R}_{m,n}^{D}$ and server-side LoRA adapters $\textbf{R}_{m,n}^{S}$ based on cut layer $c_{m,n}$.
@@
$\bullet$ Device-side FP: Device $m$ randomly draws a mini-batch of data samples $\mathcal{H}_{m,n}(t)\subseteq\mathcal{D}_{m}$ for fine-tuning the device-side pre-trained model $\textbf{w}_{m,n}^{D}$ locally. Let $\textbf{x}_{m,n}^{t}$ and $y_{m,n}^{t}$ represent the input data and its label for a mini-batch at local epoch $t$, respectively. The set of trainable LoRA adapter for the device-side pre-trained model is defined as $\textbf{R}_{m,n}^{D,t}=\{\textbf{A}_{m,n}^{1,t},\textbf{B}_{m,n}^{1,t},\dots,\textbf{A}_{m,n}^{c_{m,n},t},\textbf{B}_{m,n}^{c_{m,n},t}\}$. Device $m$ feeds a mini-batch of data into the device-side pre-trained model, and then producing the smashed data at the cut layer. The smashed data from device $m$ is denoted by
@@
$\bullet$ Smashed data transmission: Device $m$ transmits its smashed data and corresponding label to the server over wireless channels, after completing the device-side model’s FP.
@@
Stage 5 - Device-side LoRA adapters uploading. Device $m$ sends its device-side LoRA adapters $\textbf{R}_{m,n}^{D,T}$ to the server over wireless channels, after $T$ local epochs. The server updates the whole trainable LoRA adapters $\textbf{R}_{m,n}$ as follows
@@ III-A2 Transmission Delay
Here, $\varphi$ is the compression ratio for smashed data and gradient. $S(c_{m,n})$ and $\tilde{S}(c_{m,n})$ are the data size of smashed data and gradient at cut layer $c_{m,n}$, respectively. $R^{D}_{m,n}$ and $R^{S}_{m,n}$ are the transmission rates at uplink and downlink, respectively. $A(c_{m,n})$ is the data size of the device-side LoRA adapters.
@@ III-C Problem Formulation
Here, $0\leq w\leq 1$ represents the weighting factor. $D_{\min}$ and $D_{\max}$ are the minimum and maximum value of delay for device $m$ during training round $n$, respectively. Moreover $E_{\min}$ and $E_{\max}$ are the minimum and maximum value of energy consumption for device $m$ during training round $n$, respectively. The optimal problem can be formulated as follows:
$\begin{split}{\mathbf{P}_{1}:\;}&\underset{\{f_{m,n}^{S},c_{m,n}\}_{\begin{subarray}{c}m\in\mathcal{M},\\ ~{}n\in\mathcal{N}\end{subarray}}}{\text{min}}\sum_{m\in\mathcal{M}}\sum_{n\in\mathcal{N}}U(f_{m,n}^{S},c_{m,n})\\ &~{}~{}~{}~{}~{}~{}\text{s.t.}~{}c_{m,n}\in\{1,2,\cdots,I-1\},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N},\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}F_{\min}^{S}\leq f_{m,n}^{S}\leq F_{\max}^{S},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (13)
In $\mathbf{P}_{1}$, the first constraint indicates that at least one layer is on the device to reduce the risk of data leakage, and at least one layer is on the server to ensure it participates in the fine-tuning task. The secondary constraint requires that the server’s GPU frequency and average computation consumption do not surpass the upper limit $F_{\max}^{S}$, but is not less than the lower limit $F_{\min}^{S}$.
The problem $\mathbf{P}_{1}$ is a mixed integer nonlinear programming problem due to the inclusion of both continuous and integer variables. The proposed algorithm for solving this problem is discussed in the next section.
@@ IV CARD Algorithm
Considering the independence among these devices, we decompose the above problem $\mathbf{P}_{1}$ into multiple sub-problems to allow for efficient solutions. Each sub-problem is to jointly determine the optimal cut layer and computing resource allocation for device $m$ at training round $n$, which can be expressed as follows:
$\begin{split}{\mathbf{P}_{2}:\;}&\underset{f_{m,n}^{S},c_{m,n}}{\text{min}}~{}U(f_{m,n}^{S},c_{m,n})\\ &~{}\text{s.t.}~{}c_{m,n}\in\{1,2,\cdots,I-1\},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N},\\ &~{}~{}~{}~{}~{}F_{\min}^{S}\leq f_{m,n}^{S}\leq F_{\max}^{S},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (14)
IV-1 Upper-Layer Problem
$\begin{split}{\mathbf{P}_{3}:\;}&\underset{f_{m,n}^{S}}{\text{min}}~{}U(f_{m,n}^{S})\\ &\text{s.t.}~{}F_{\min}^{S}\leq f_{m,n}^{S}\leq F_{\max}^{S},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (15)
$(f_{m,n}^{S})^{\star}=\begin{cases}F_{\min}^{S},\hfill~{}~{}\text{if}~{}Q<F_{\min}^{S},\\ Q,\hfill~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}F_{\min}^{S}\leq Q\leq F_{\max}^{S},\\ F_{\max}^{S},\hfill~{}~{}\text{if}~{}F_{\max}^{S}<Q,\end{cases}$ (16)
IV-2 Lower-Layer Problem
$\begin{split}{\mathbf{P}_{4}:\;}&\underset{c_{m,n}}{\text{min}}~{}U(c_{m,n})\\ &\text{s.t.}~{}c_{m,n}\in\{1,2,\cdots,I-1\},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (17)
The function in the problem $\mathbf{P}_{4}$ is non-convex because the data sizes (e.g., device-side LoRA adapter, smashed data, and its gradient) are arbitrary functions related to the cut layer. The optimal cut layer can be found by the brute-force approach, due to the finite number of transformer layers in an LLM. The CARD algorithm is summarized in Alg. 1.

### v2

@@ I Introduction
Parameter efficient fine-tuning (PEFT) has been explored to reduce trainable parameters in LLMs, enhancing task-specific performance in edge networks [3]. By fine-tuning a small subset of parameters, PEFT achieves performance comparable to full-model fine-tuning. In particular, as a popular PEFT method, low-rank adaptation (LoRA) keeps the pre-trained model weights fixed and introduces trainable low-rank matrices within each transformer layer. In edge networks, many private institutions or companies face limitations in accessing the data stored on edge devices due to concerns about data privacy and security. This issue is exacerbated when fine-tuning requires the transfer of data to centralized servers, raising the risk of exposure to malicious actors [4].
@@
Federated learning (FL) offers a promising solution to privacy concerns by enabling decentralized fine-tuning directly on multiple owners, sharing only model updates instead of raw data [5]. However, applying FL for fine-tuning LLMs in edge networks remains challenging, as the process of fine-tuning the entire LLM is performed on the participating devices. This approach requires enormous computational overhead and memory, which hinders its feasibility on edge devices [6]. For example, fine-tuning a T5-Large model with LoRA requires memory of 7.1 GB in a mini-batch training, while a Jetson Nano, an edge AI platform, has only 4 GB of RAM [7].
@@
To address this issue, split learning (SL) offers an alternative to reduce the computational burden on edge devices while preserving data privacy [8]. In SL, the initial layers of a model are processed locally on the device, and only the intermediate activations (not the raw data) are sent to the server, which then completes the remaining layers of the model [9]. In [10], an SL-based fine-tuning approach is introduced, where the devices fine-tune the initial layers while servers handle the rest. Similarly, a split LoRA-based fine-tuning framework is presented, where the embedding and task modules are executed on devices, with the encoder kept on the server [11]. Most existing literature focuses on static split strategies and neglects server-side energy consumption.
@@
However, optimizing LLMs in edge networks is challenging in achieving the best possible latency and energy efficiency performance. Firstly, using a static cut is suboptimal due to heterogeneous devices and the dynamic channel in edge networks. Secondly, static server resource configurations can lead to inefficient energy use as workloads vary with the cut layer. Therefore, an adaptive strategy that dynamically adjusts the cut layer and server’s computational resource based on real-time conditions is necessary to optimize both latency and energy efficiency.
@@
In this letter, firstly, we propose an energy-efficient SL framework for fine-tuning LLMs, which aims to enhance training and energy efficiency in edge networks. Secondly, we analyze the factors of the training delay and energy consumption in detail, then formulate an optimization problem to minimize both the training delay and the server’s energy consumption. Thirdly, to solve this problem, we design an algorithm of Cut lAyer and computing Resource Decision, named CARD. It decomposes this problem into independent sub-problems and then jointly determines the cut layer selection and server’s computation resource allocation. Finally, extensive simulations demonstrate the performance of the proposed approach through a physical platform with 5 NVIDIA Jetson edge devices. The main contributions of this work are summarized as follows:
@@ II-A LoRA-Based LLM Fine-Tuning
A typical LoRA-based LLM fine-tuning scenario is considered, in which multiple edge devices collaborate with an access point (AP) to fine-tune a shared pre-trained LLM. Each device trains only the initial layers of the LLM, reducing local computation and memory requirements. The AP with an edge server coordinates to complete the remaining training process with each device, as depicted in Fig. 1. This approach is ideal for resource-constrained environments, enabling collaborative fine-tuning while enhancing training and energy efficiency.
Let $\textbf{W}\in\mathbb{R}^{P\times Q}$ denote a pre-trained weight matrix of the LLM, with its corresponding model update expressed as $\textbf{W}+\Delta\textbf{W}=\textbf{W}+\textbf{A}\textbf{B}$. Here, $\textbf{A}\in\mathbb{R}^{P\times Z}$ and $\textbf{B}\in\mathbb{R}^{Z\times Q}$ are a pair of decomposed low-rank matrices, and $Z\ll\min(P,Q)$, as shown in Fig. 2. In the LoRA-based fine-tuning process, the pair of decomposed low-rank matrices are used for adaptation, while the weight matrix W is frozen. The edge device is indexed by $m\in\mathcal{M}=\{1,2,...,M\}$, and its local dataset is denoted by $\mathcal{D}_{m}=\{\textbf{x}_{m,k},y_{m,k}\}_{k=1}^{D_{m}}$. Here, $\textbf{x}_{m,k}$ and $y_{m,k}$ are the $k$-th input data and its corresponding label. A device-side pre-trained model and its associated set of trainable LoRA adapters are denoted by $\textbf{W}_{m}^{D}$ and $\textbf{R}_{m}^{D}$, respectively. $\textbf{R}_{m}^{D}=\{\textbf{A}_{m}^{1},\textbf{B}_{m}^{1},\dots,\textbf{A}_{m}^{c_{m}},\textbf{B}_{m}^{c_{m}}\}$, where $\textbf{A}_{m}^{i}$ and $\textbf{B}_{m}^{i}$ are the pair of low-rank matrices of $i$-th LoRA adapter, and $c_{m}$ is the cut layer.
@@
For the edge server deployed in the AP, a server-side pre-trained model and its corresponding set of trainable LoRA adapters are denoted by $\textbf{W}_{m}^{S}$ and $\textbf{R}_{m}^{S}$, respectively. Similarly, $\textbf{R}_{m}^{S}=\{\textbf{A}_{m}^{c_{m}+1},\textbf{B}_{m}^{c_{m}+1},\dots,\textbf{A}_{m}^{I},\textbf{B}_{m}^{I}\}$, where $I$ is the total number of transform layers in the LLM. The global pre-trained LLM is denoted by $\textbf{W}=\{\textbf{W}_{m}^{D};\textbf{W}_{m}^{S}\},~{}\forall m\in\mathcal{M}$. The objective of split LoRA-based fine-tuning is to find the optimal LoRA adapter configuration that minimizes the global loss across all participating devices, i.e.,
@@ II-B Proposed SL Framework for fine-tuning LLMs
In this section, the proposed SL framework for fine-tuning LLMs is presented in detail. The basic idea is to leverage the SL method, allowing devices to fine-tune the first layers of the LLM while the server handles the remaining layers. Before the initiation of fine-tuning, the parameters in LoRA adapters are initialized randomly. After that, the LLM is executed between devices and the server in consecutive training rounds until the optimal LoRA adapters are identified. The training round is indexed by $n\in\mathcal{N}=\{1,2,...,N\}$, and the detailed workflow of the proposed framework is as follows.
Stage 1 - LLM splitting. The edge server selects device $m$ to execute LLM fine-tuning collectively. Then, it splits the trainable LoRA adapters into device-side LoRA adapters $\textbf{R}_{m,n}^{D}$ and server-side LoRA adapters $\textbf{R}_{m,n}^{S}$ based on cut layer $c_{m,n}$.
@@
$\bullet$ Device-side FP: Device $m$ randomly draws a mini-batch of data samples $\mathcal{H}_{m,n}(t)\subseteq\mathcal{D}_{m}$ for fine-tuning the device-side pre-trained model $\textbf{w}_{m,n}^{D}$ locally. Let $\textbf{x}_{m,n}^{t}$ and $y_{m,n}^{t}$ represent the input data and its label for a mini-batch at local epoch $t$, respectively. The set of trainable LoRA adapter for the device-side pre-trained model is defined as $\textbf{R}_{m,n}^{D,t}=\{\textbf{A}_{m,n}^{1,t},\textbf{B}_{m,n}^{1,t},\dots,\textbf{A}_{m,n}^{c_{m,n},t},\textbf{B}_{m,n}^{c_{m,n},t}\}$. Device $m$ feeds a mini-batch of data into the device-side pre-trained model, producing the smashed data at the cut layer. The smashed data from device $m$ is denoted by
@@
$\bullet$ Smashed data transmission: Device $m$ transmits its smashed data and corresponding label to the server over wireless channels after completing the device-side model’s FP.
@@
Stage 5 - Device-side LoRA adapters uploading. Device $m$ sends its device-side LoRA adapters $\textbf{R}_{m,n}^{D,T}$ to the server over wireless channels after $T$ local epochs. The server updates the whole trainable LoRA adapters $\textbf{R}_{m,n}$ as follows
@@ III-A2 Transmission Delay
Here, $\varphi$ is the compression ratio for smashed data and gradient. $S(c_{m,n})$ and $\tilde{S}(c_{m,n})$ are the data size of smashed data and gradient at cut layer $c_{m,n}$, respectively. $R^{D}_{m,n}$ and $R^{S}_{m,n}$ are the transmission rates at uplink and downlink, respectively. The transmission rate is converted from signal-to-noise ratio (SNR) by the channel quality indication (CQI) to modulation and coding scheme (MCS) mapping table [12], i.e., $R_{m,n}=B_{m,n}y(\text{SNR}_{m,n})$. Here, $B_{m,n}$ is the bandwidth allocated for the $m$-th device, and $y(\cdot)$ is the conversion function mapping SNR to the spectral efficiency. $A(c_{m,n})$ is the data size of the device-side LoRA adapters.
@@ III-C Problem Formulation
Here, $0\leq w\leq 1$ represents the weighting factor. $D_{\min}$ and $D_{\max}$ are the minimum and maximum value of delay for device $m$ during training round $n$, respectively. Moreover, $E_{\min}$ and $E_{\max}$ are the minimum and maximum value of energy consumption for device $m$ during training round $n$, respectively. $D_{\max}$ and $E_{\min}$ can be obtained when $c_{m,n}=I$ and $f_{m,n}^{S}=F_{\min}^{m,S}$. $D_{\min}$ and $E_{\max}$ can be obtained when $c_{m,n}=0$ and $f_{m,n}^{S}=F_{\max}^{S}$. The optimal problem can be formulated as:
$\begin{split}{\mathbf{P}_{1}:\;}&\underset{\{f_{m,n}^{S},c_{m,n}\}_{\begin{subarray}{c}m\in\mathcal{M},\\ ~{}n\in\mathcal{N}\end{subarray}}}{\text{min}}\sum_{m\in\mathcal{M}}\sum_{n\in\mathcal{N}}U(f_{m,n}^{S},c_{m,n})\\ &~{}~{}~{}~{}~{}~{}\text{s.t.}~{}c_{m,n}\in\{0,1,\cdots,I\},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N},\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}F_{\min}^{m,S}\leq f_{m,n}^{S}\leq F_{\max}^{S},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (13)
Here, the first constraint guarantees the feasibility of the cut layer selection. The secondary constraint requires that the server’s GPU frequency does not surpass the upper limit $F_{\max}^{S}$, but is not less than the lower limit $F_{\min}^{m,S}$. Here, $F_{\max}^{S}$ is the maximum frequency the server’s can reach, and $F_{\min}^{m,S}=f_{m}^{D}\delta_{m}^{D}\sigma_{m}^{D}/(\delta^{S}\sigma^{S}),\forall m$, which stems from the assumption that the computational power of the server is more powerful than that of devices to ensure the effectiveness of split learning.
The problem $\mathbf{P}_{1}$ is a mixed integer nonlinear programming problem due to the inclusion of continuous and integer variables. The proposed algorithm for solving this problem is discussed in the next section.
@@ IV CARD Algorithm
Considering the independence among these devices, we decompose problem $\mathbf{P}_{1}$ into multiple sub-problems to allow for efficient solutions. Each sub-problem is to jointly determine the optimal cut layer and computing resource allocation for device $m$ at training round $n$, which can be expressed as follows:
$\begin{split}{\mathbf{P}_{2}:\;}&\underset{f_{m,n}^{S},c_{m,n}}{\text{min}}~{}U(f_{m,n}^{S},c_{m,n})\\ &~{}\text{s.t.}~{}c_{m,n}\in\{0,1,\cdots,I\},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N},\\ &~{}~{}~{}~{}~{}F_{\min}^{m,S}\leq f_{m,n}^{S}\leq F_{\max}^{S},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (14)
IV-1 Upper-Layer Problem
$\begin{split}{\mathbf{P}_{3}:\;}&\underset{f_{m,n}^{S}}{\text{min}}~{}U(f_{m,n}^{S})\\ &\text{s.t.}~{}F_{\min}^{m,S}\leq f_{m,n}^{S}\leq F_{\max}^{S},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (15)
$(f_{m,n}^{S})^{\star}=\begin{cases}F_{\min}^{m,S},\hfill~{}~{}\text{if}~{}Q<F_{\min}^{m,S},\\ Q,\hfill~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}F_{\min}^{m,S}\leq Q\leq F_{\max}^{S},\\ F_{\max}^{S},\hfill~{}~{}\text{if}~{}F_{\max}^{S}<Q,\end{cases}$ (16)
IV-2 Lower-Layer Problem
$\begin{split}{\mathbf{P}_{4}:\;}&\underset{c_{m,n}}{\text{min}}~{}U(c_{m,n})\\ &\text{s.t.}~{}c_{m,n}\in\{0,1,\cdots,I\},~{}\forall m,~{}n\in\mathcal{M},~{}\mathcal{N}.\end{split}$ (17)
The function in the problem $\mathbf{P}_{4}$ is non-convex because the data sizes (e.g., device-side LoRA adapter, smashed data, and its gradient) are arbitrary functions related to the cut layer. Due to the finite number of transformer layers in an LLM, the optimal cut layer can be found by the brute-force approach. The CARD algorithm is summarized in Alg. 1. Its time complexity is $O(I)$, which requires $I+1$ iterations to find the optimal cut layer and the server’s GPU frequency.